SYNCHRONOUSLY SCANNED OPO CARS: A NEW TOOL FOR MOLECULAR SPECTROSCOPY

\underline{PETER C. CHEN}, Spelman College, 350 Spelman Ln. Box 307, Atlanta, Georgia, 30314.

Results are presented from the development of a new nonlinear spectroscopic method for molecular vibrational spectroscopy. This method uses the Optical Parametric Oscillator (OPO) and synchronous scanning to overcome problems that have limited the scan range for nonlinear Raman spectroscopy in the past. This new improvement transforms nonlinear Raman spectroscopy into a more appealing tool for studying molecular species.